### v1

# Multilingual Synopses of Movie Narratives: A Dataset for Story Understanding

Abstract
Story video-text alignment, a core task in computational story understanding, aims to align video clips with corresponding sentences in their descriptions. However, progress on the task has been held back by the scarcity of manually annotated video-text correspondence and the heavy concentration on English narrations of Hollywood movies. To address these issues, in this paper, we construct a large-scale multilingual video story dataset named Multilingual Synopses of Movie Narratives (M-SyMoN), containing 13,166 movie summary videos from 7 languages, as well as manual annotation of fine-grained video-text correspondences for 101.5 hours of video. Training on the human annotated data from M-SyMoN outperforms the SOTA methods by 15.7 and 16.2 percentage points on Clip Accuracy and Sentence IoU scores, respectively, demonstrating the effectiveness of the annotations. As benchmarks for future research, we create 6 baseline approaches with different multilingual training strategies, compare their performance in both intra-lingual and cross-lingual setups, exemplifying the challenges of multilingual video-text alignment.
Multilingual Synopses of Movie Narratives: A Dataset for Story Understanding
1 Introduction
Computational story understanding aims to empower AI systems with the ability to delve into the intricacies of diverse stories, unlocking their deep semantics such as character motivations and intentions Emelin et al. (2021); Rashkin et al. (2018), event structures Chambers and Jurafsky (2008); Du et al. (2021), and social relationships among story characters Elson et al. (2010); Chaturvedi et al. (2016); Kim and Klinger (2019). In recent years, computation story understanding has garnered significant research interest Dong et al. (2023); Sang et al. (2022); Andrus et al. (2022); Xu et al. (2022); Han et al. (2023) and many story understanding tasks Wu and Krahenbuhl (2021); Choi et al. (2021) have emerged.
Story video-text alignment is a fundamental task of computational story understanding, which aims to find the best correspondence between a sequence of video clips and a sequence of sentences (see an alignment example in Figure 1). Different from traditional video-text retrieval that relies on keyword or temporal cue matching Wang et al. (2021b), story video-text alignment requires various story understanding abilities such as causal chain reasoning, mental state description, and long-range context understanding Sun et al. (2022). Establishing such video-text correspondence will facilitate many applications such as text-to-video generation Liu et al. (2019); Balaji et al. (2019), and visual story generation Huang et al. (2019). Thus, the story video-text alignment task has recently attracted increasing attention Dogan et al. (2018); Wang et al. (2021a); Sun et al. (2024).
Figure 1: An example alignment between a video clip sequence and a sentence sequence from M-SyMoN. One text chuck may correspond to several video clips, while some video clips may not match any textual description.
However, a major obstacle of this task stems from the scarcity of annotated data, as it is costly and time-consuming to manually annotate sentence-level alignments between text and clip sequences. The story understanding datasets used in previous studies are limited for several reasons. First, although the Large Scale Movie Description Challenge (LSMDC) dataset Rohrbach et al. (2017) manually aligns 158 hours of movies and audio descriptions, it only considers one-to-one matching between movie clips and audio descriptions and its audio descriptions are designed for visually impaired audiences, containing excessive details. Second, the YouTube Movie Summary (YMS) dataset Dogan et al. (2018) is small and only contains 6.7 hours of video and texts, and thus is typically employed for video-text alignment evaluation. Lastly, Hollywood movies and English narrations are dominant in existing story understanding datasets Huang et al. (2020); Soldan et al. (2022); Sun et al. (2022); Lu et al. (2023), neglecting the importance of language and cultural diversity.
To address the aforementioned limitations, we construct a multilingual video story dataset named Multilingual Synopses of Movie Narratives (M-SyMoN). M-SyMoN is sourced from movie recap videos on YouTube and contains 13,166 videos spanning 7 languages and totalling 2,136 hours. Furthermore, we manually annotated a portion of this dataset, providing exact video-text correspondence for 480 videos or 101.5 video hours. Compared to LSMDC and YMS, our annotated video-text alignment subset is large-scale, contains one-to-many matching and unmatched text or video clips, and covers 7 different languages.
We further investigate the multi-lingual characteristics of the dataset, and make the following observations. First, in the intra-lingual setup, compared to translating all languages to English for training and inference, additional language-specific fine-tuning on weakly supervised data brings an average improvement of 5.4 percentage points across 7 languages. Adding a small portion of the manual annotations further boosts performance. Second, in the cross-lingual setup, for source and target languages that are linguistically similar (e.g., Spanish to Portuguese or French), the transfer performance is generally good; for source and target languages that are different (e.g., Chinese to English or Hindi), the transfer performance is quite limited. Third, an out-of-domain evaluation on the YMS dataset shows that training on the weakly supervised data from M-SyMoN outperforms the state of the art methods by 12.3 and 13.2 percentage points on Clip Accuracy and Sentence IoU scores. Moreover, adding manually annotated video-text alignment data further improves the performance by 2.4 and 3.0 percentage points, indicating the utility of our annotated alignment data.
The main contributions of this work are summarized as follows:
  • We construct a large-scale multilingual video story understanding dataset named M-SyMoN, containing 13,166 videos in 7 languages and totaling 2,136 hours.
  • We manually annotate the fine-grained alignment between video clips and sentences of 480 videos for a total of 101.5 hours. [^1] The video URLs and video-text alignment annotations in M-SyMoN will be made publicly available.
  • We create a number of multilingual video-text alignment methods to benchmark the M-SyMoN dataset. Extensive results on both M-SyMoN and YMS demonstrate the significance of our multilingual dataset and the utility of the human video-text alignment annotations.
2 Related Work
2.1 Story Video-Text Alignment
Story video-text alignment involves aligning sequences of video clips, typically from movies, with text captions Cour et al. (2008); Tapaswi et al. (2015); Wang et al. (2021a); Dogan et al. (2018). The common approach is to first learn a video-text similarity metric and then calculate the alignment using dynamic programming (e.g., DTW) Zhang et al. (2023b); Dvornik et al. (2021). Recent methods encode videos with 3D-convolution Sun et al. (2024) or ViT Li et al. (2023) and texts with Transformer models Han et al. (2022); Li et al. (2023); Zhang et al. (2023a), then optimize alignment via soft-DTW Zhang et al. (2023a); Han et al. (2022) or calculate the alignment from video-text similarity Zhang et al. (2023b).
Due to the lack of sentence-level video-text alignment annotations, most models are trained with timestamp-based weak supervision. Although the LSMDC dataset Rohrbach et al. (2017) contains a large amount of manual annotation, its overly fine-grained text data interferes with event understanding Sun et al. (2022); it also lacks one-to-many matching and unmatched items. The YMS dataset Dogan et al. (2018) contains only 6.7 hours of movie summaries with exact video-text alignment and cannot serve as a comprehensive test benchmark.
2.2 Movie Story Datasets
Movies are a popular source of video-text story content. LSMDC Rohrbach et al. (2017) and Movie Audio Description (MAD) Soldan et al. (2022) provide movie clips with audio descriptions for the visually impaired. Although the audio descriptions are accurate, they deviate significantly from realistic styles of story narration. The Condensed Movie Descriptions (CMD) dataset Bain et al. (2020) offers 7 to 11 key clips per movie with one-sentence descriptions. The Pororo dataset Kim et al. (2017) includes 20-minute cartoon episodes with in-show conversations and human-written descriptions. Although CMD and Pororo captions match the clips, they may not form complete storylines. The CVSV Lu et al. (2023), YMS Dogan et al. (2018), and SyMoN Sun et al. (2022) datasets collect YouTube movie summaries, similar to M-SyMoN, but CVSV and SyMoN lack human annotations.
Although there are many story understanding dataset, multilingual video story datasets are scarce. To our knowledge, the only dataset is Movie101v2 Yue et al. (2024), which contains 46K Chinese video-caption pairs from 203 movies, where the Chinese text are translated to English using GPT-3.5-turbo. The dataset also lacks human-annotated video-text correspondence.
In comparison, M-SyMoN is large-scale and multilingual, containing movie summary videos in 7 languages and 101.5 hours of video with exact video-text correspondences.
2.3 Multilingual Story Understanding
There are several tasks related to multilingual story understanding, including event-causal inference Lai et al. (2022), story question answering Ateeq et al. (2023), story tag classification Tikhonov et al. (2021), and story generation Razumovskaia et al. (2024). However, most of them merely focus on text rather than video stories.
Methods for multilingual story understanding mainly fall into two categories: (1) translating training/test data to English Ponti et al. (2020) or prompts to the target language Lin et al. (2022), and (2) directly fine-tuning on non-English text using a multilingual Pretrained Language Model (PLM) Tikhonov et al. (2021) or a PLM for the target language Lai et al. (2022). In this paper, we benchmark M-SyMoN on multilingual story video-text alignment with both types of methods.
3 Task Formulation
Given a consecutive sequence of video clips $V=(v_{1},v_{2},...,v_{m})$ and a consecutive sequence of sentences $T=(t_{1},t_{2},..t_{n})$, the video-text alignment task aims to find the alignment between these two sequences by learning a function $f$ that maps each input sentence $t_{i}$ to its corresponding video clips:
$\mathcal{P}=\{\dots,(t_{i},v_{f(i)}),\dots\}$ (1)
where $f(i)$ refers to the indexes of the video clip aligned with the sentence $t_{i}$. Note that $f(i)$ can be None, a single index, or an index sequence, denoting that there is no video clip, one video clip, or multiple video clips aligned with $t_{i}$, respectively.
Languages English Chinese Spanish French Portuguese Hindi Russian
Video Count 5,193 2,683 1,595 1,193 1,070 749 683
Video Hours 869 390 285 102 209 173 108
Vocabulary Size 40,116 6,269 50,050 33,967 40,676 20,896 64,827
Average Video Length (minutes) 9.5 8.7 10.7 5.1 11.7 13.9 9.5
Average number of Sentences in a Narration 131 78.8 82.6 138 134 126 151
Average number of words in a Narration 1,717 3,234 1,923 2,226 2,737 2,656 1,623
Movie Count 2,440 1,236 1,217 877 811 462 542
Number of Annotated Videos 98 31 75 84 57 63 72
Annotated Video Hours 23.42 7.23 14.17 14.17 14.17 14.17 14.17
Table 1: The statistics of M-SyMoN. For Chinese, we regard the number of unique characters as its vocabulary size.
4 Dataset Construction
In this section, we describe the details of constructing the M-SyMoN dataset.
4.1 Data Source and Statistics
For data collection, we first identify YouTube channels with movie recap videos in the target languages. Keywords such as “movie summary in <language>” and “movie recap in <language>” are used to search for these channels. We then download all videos and their accompanying subtitles from the identified channels. Videos without subtitles and that are not movie summaries are discarded.
This yields a total of 13,166 videos or 2,136 video hours in 7 languages, including English, Chinese, Spanish, French, Portuguese, Hindi, and Russian. We list the statistics in Table 1. M-SyMoN contains summaries for 5,960 movies and TV shows, of which 1,515 have more than one summary.
4.2 Human Annotation
We hire a professional team of annotators with rich translation experience from Flitto [^2] https://www.flitto.com to annotate exact video-text correspondence in M-SyMoN. In total, we annotated 480 videos spanning 101.5 hours. The amount of annotated videos in each language are in the last two lines of Table 1.
Before human annotation, we automatically divided the video subtitle text into sentences. For each sentence, we ask the annotators to locate the start and end times of the video segment described by the sentence. If a sentence is not grounded in the video, it is marked as “unmatched”. We removed the audio from the videos to eliminate shortcut features for alignment from the audio. Note that human-annotated video-text correspondence is different from temporal correspondence, because the narration in narrated movie summary videos is sometimes faster or slower than the video, and some narrator may add commentary not grounded in the video.
Annotation Quality. To evaluate annotation quality, we also employ a student who is familiar with the task to annotate a small validation set for each language. The annotation quality is evaluated as the IoU between the durations annotated by the student and the durations annotated by the annotators. If the IoU for a particular language falls below 60%, the annotators were asked to redo the annotations until the IoU reaches 60%. The final average IoU across 7 languages is 83.1%, indicating substantial agreement between the annotators. See Appendix B for annotation details.
4.3 Data Preprocessing
Multilingual Punctuation Restoration. We acquire text descriptions directly from YouTube subtitles. In some cases, the text descriptions are derived from automatic speech recognition tools and are unpunctuated. As punctuated texts are required for downstream tasks, we train a Transformer-based model to restore punctuations in all unpunctuated text. We do not use off-the-shelf punctuation restoration models because most models do not include all the languages in M-SyMoN Chordia (2021); Frank and Böhme (2021). See Appendix C for details on the punctuation restoration model.
Scene Segmentation. We divide the videos into clips using TransNet V2 Souček and Lokoč (2020) which detects hard camera cuts. Each clip, containing a continuous shot between two hard camera cuts, is roughly 2.4 seconds long.
Offensive Language Filtering. To the best of our abilities, we created a list of offensive terms and filtered out videos containing those phrases.
4.4 Dataset Split
We divide M-SyMoN into two parts: (1) the human annotated portion: 480 videos from 7 languages manually annotated with exact video-text correspondence; (2) the weakly supervised portion: The entire M-SyMoN dataset after removing all movies that appear in the annotated portion (see Appendix A.1 for details).
For the weakly supervised data, each video clip is matched to the text segment that spans its duration, producing a rough video-text correspondence regarded as the weakly supervised training set.
For the human annotated data, we further split it into a training set (20%), a validation set (20%), and a test set (60%), with the portions distributed evenly among 7 languages. Table 2 shows the number of clip-sentence pairs in each split.
Languages English Chinese Spanish French Portuguese Hindi Russian
Weakly supervised training set 730,112 365,312 166,530 60,676 127,788 90,560 84,556
Supervised training set 6,649 1,860 5,989 5,799 4,205 7,152 5,789
Validation set 7,201 1,616 4,066 4,717 6,865 7,603 8,474
Test set 16,208 5,441 15,185 12,560 15,356 19,652 14,664
Table 2: The number of clip-sentence pairs in each data split.
4.5 Dataset Analysis
We hypothesize that the choices of movies in each language may have cultural characteristics. Due to space limitations, we analyze the country of release and themes of the movies in this section. Additional statistics of the genre, year of release, language of release, and narrative structures are illustrated in Appendix A.2.
As most movie metadata are not available from the movie recap videos, we use ChatGPT (gpt-3.5-turbo-0125) to acquire the meta information. Specifically, we input the video title, the video-level description featured on YouTube, and first 5 sentences of the movie narration into ChatGPT and ask ChatGPT to output the movie title and metadata.
Figure 2: The top 5 countries of release for movies recapped in each language.
Figure 2 reveals interesting correlation between the language and the origin of the movie. The movie summaries in each language are typically about domestic movies (i.e., movies from countries where the language is widely used) or from the United States, reflecting the influence of Hollywood. For instance, 22.5% of Portuguese videos are about Brazilian movies. Interestingly, large percentage of videos in Chinese (46.9% from China Mainland, 7.6% from Hong Kong) and Russian (44.0%) describe domestic movies than other languages. In comparison, on average 20.4% of videos in other non-english languages describe domestic movies.
Figure 3 reveals variations in thematic preferences across languages. While themes like “Friendship” and “Family” appear universally popular, others exhibit uneven distributions. For instance, 2.7% (vs. 0.7% in other languages on average) of Russian movies have the theme "Isolation". The theme “Identity” accounts for only 1.8% Chinese videos (vs. an average of 4.6% in all other languages). These differences highlight the importance of representation of multiple cultures and languages in story understanding datasets.
Figure 3: The top 5 themes for movies recapped in each language.
5 Methodology
This section first details our base model for video-text alignment and then introduces its several variants with different multilingual training strategies.
5.1 Base Model
To achieve the sentence-level alignments between text and clip sequences, we adopt a three-stage framework widely used in recent studies Dvornik et al. (2021); Zhang et al. (2023b). Specifically, our model first employs video and text encoders to obtain the video clip representation and the text representation as follows.
Video Encoder. Given a video with $M$ clips, we represent each clip by randomly selecting three frames from it and dividing each frame into $H\times W$ patches. We then utilize a trainable MLP projection layer to map these patches into video tokens, which are fed into the video encoder Swin Transformer Liu et al. (2021) to obtain the visual representation of each frame. An average pooling layer is applied over the three frames of each video clip to obtain the video clip representation.
Text Encoder. Because our goal is to perform multilingual video-text alignment in this paper, we adopt a pre-trained cross-lingual cross-modal model named CCLM Zeng et al. (2023) to initialize the parameters in Swin Transformer and the text encoder. Following CCLM, we employ a pre-trained multi-layer Transformer, i.e., XLM-R Conneau et al. (2020) as the text encoder. We then regard the last hidden representation of the [CLS] token as the representation of each sentence. Formally, let $\bm{t}_{i}$ and $\bm{v}_{i}$ denote the encoded features for the $i^{\text{th}}$ sentence and the $i^{\text{th}}$ video clip.
For every clip-text pair, we randomly sample $K$ hard negatives from the same video. With the sampled negative text features and video features, we fine-tune the encoders by minimizing the contrastive InfoNCELoss (Oord et al., 2018) below:
$\mathcal{L}_{\text{InfoNCE}}=$ $\frac{1}{N}\sum_{i=1}^{N}\left[-\log\left(\frac{\exp(\bm{v}_{i}^{\top}\bm{t}_{i}/\tau)}{\sum_{j=1}^{K}\exp(\bm{v}_{i}^{\top}\bm{t}_{j}/\tau)}\right)\right.$ (2)
$\left.-\log\left(\frac{\exp(\bm{v}_{i}^{\top}\bm{t}_{i}/\tau)}{\sum_{j=1}^{K}\exp(\bm{v}_{j}^{\top}\bm{t}_{i}/\tau)}\right)\right]$
where $N$ is the total number of training samples and $K$ is the number of negative samples.
During inference, we first obtain the representation of each clip and sentence, and then calculate the cosine similarity between every clip-sentence pair. Finally, we find the global clip-sentence alignment from the similarity scores of each pair by resorting to a sequence alignment algorithm named Double Drop Dynamic Time Warp (Drop-DTW) Dvornik et al. (2021), detailed in Appendix D.1.
English Chinese Spanish French Portuguese Hindi Russian Average
Weakly supervised
CCLM-multilingual 9.3 16.3 11.8 8.8 9.9 5.8 9.3 10.2
CCLM-individual 26.6 29.3 17.4 16.6 16.3 10.1 14.6 18.7
CCLM-translate 22.2 20.1 16.0 17.5 14.1 11.2 13.8 16.4
CCLM-two-stage 26.9 37.7 19.1 20.2 18.7 13.1 17.0 21.8
Supervised
CCLM-translate-supervision 24.3 20.1 17.8 19.7 16.0 12.6 15.1 17.9
CCLM-two-stage-supervision 27.7 38.9 19.8 21.0 19.5 12.9 17.5 22.5
Table 3: Intra-lingual results based on F1 score. The “Average” is the mean of 7 languages. The highest number in each column is in bold, the second highest is in bold italic.
5.2 Multilingual Fine-Tuning Methods
With the base model, we benchmark the M-SyMoN dataset by establishing 4 weakly supervised methods, which share the same model architecture but are trained on the entire weakly supervised training set of M-SyMoN in Section 4.4 with different training strategies:
  • Multilingual training (CCLM-multilingual). We combine the weakly supervised training sets of all 7 languages and train a unified multilingual model on the combined dataset.
  • Individual training (CCLM-individual). Using weakly supervised training data, we fine-tune the base model on each language and obtain 7 language-specific models.
  • Translational training (CCLM-translate). We first translate all the weakly supervised training data and the test data in M-SyMoN to English with off-the-shelf translation model NLLB-200 Costa-jussà et al. (2022). We then fine-tune the base model on the translated data to obtain a unified model.
  • Two-stage training (CCLM-two-stage). On top of the trained CCLM-translate model, we further fine-tune on each language separately to obtain 7 language-specific models.
To show the usefulness of our manually annotated video-text alignment dataset, we further fine-tune two aforementioned methods on the human annotated training set in Section 4.4:
  • CCLM-translate-supervision. We translate the non-English languages in the human annotated data to English and fine-tune the CCLM-translate model on the translated data.
  • CCLM-two-stage-supervision. We fine-tune the CCLM-two-stage model on the human annotated data of each language separately.
6 Experiments
In this section, we conduct extensive experiments in intra-lingual, cross-lingual, and out-of-domain setups to show the usefulness of M-SyMoN.
6.1 Experimental Setup
For CCLM-translate and CCLM-multilingual, we initialize their parameters from the pre-trained CCLM-base model Zeng et al. (2023) and fine-tune them for 20 epochs with an initial learning rate of $4\times 10^{-5}$ and cosine learning rate decay. For CCLM-two-stage, we initialize its parameters from the CCLM-translate model and fine-tune it for 20 epochs with an initial learning rate of $4\times 10^{-6}$ and cosine learning rate decay. Random augmentation and weight decay of 0.2 are applied.
6.2 Evaluation Metrics
Following Dogan et al. (2018), we use two evaluation metrics. Clip accuracy is defined as the temporal proportion of correctly aligned video segments. Sentence IoU is defined as the intersection-over-union between the aligned video durations and the groundtruth durations. Due to space limitations, we mainly report the F1 score, i.e. the harmonic mean between Clip Accuracy and Sentence IoU in this section and defer detailed results to Appendix D.2.
6.3 Intra-Lingual Results
We report the intra-lingual results in Table 3. Here, each method is trained and evaluated on data in the same language.
First, we find that training the base model on all languages together performs significantly worse than training on each language separately. Specifically, CCLM-individual outperforms CCLM-multilingual by 4.3-17.3 percentage points. This suggests that language-specific features are important for multilingual video-text alignment. Second, the CCLM-translate baseline is on par with CCLM-individual and significantly outperforms CCLM-multilingual. Moreover, the two-stage baseline (i.e., CCLM-two-stage) leverages the benefits of pre-training on a large multilingual dataset and subsequent fine-tuning on a specific language dataset, which outperforms CCLM-individual and CCLM-translate by 3.1 and 5.4 percentage points on average. Finally, after finetuning with our manually annotated data, it yields consistent improvements on top of both CCLM-translate and CCLM-two-stage models across the 7 languages, and CCLM-two-stage-supervision achieves the best average performance among all the compared methods. These observations demonstrate the usefulness of the manually annotated video-text alignment data.
Figure 4: Cross-lingual transfer results of CCLM-two-stage based on F1 score. The language names are abbreviated as: English =“en”, Chinese = “ch”, Spanish = “es”, French = “fr”, Portuguese “pt”, Hindi = “hi”, Russian = “ru”.
Figure 5: Case Study. The original movie descriptions are in Chinese. The CCLM-translate baseline is trained and evaluated on the text translated to English with NLLB-200 Costa-jussà et al. (2022). The CCLM-two-stage baseline is evaluated on the Chinese text, and we show the human-translated English text here for clear illustration. Green boxes mark correct alignments and red boxes mark incorrect alignments.
6.4 Cross-Lingual Transfer Results
We employ the CCLM-two-stage model to evaluate the cross-lingual performance of every source-target language pair. The results are reported in Figure 4.
First, the highest value of each column appears on the diagonal. This is intuitive because the model is trained and evaluated on the same language. Second, linguistically similar languages transfer well to each other. For example, related languages like Spanish and Portuguese, generally obtain good cross-lingual performance. This likely stems from shared vocabulary, sub-word tokens, and grammatical structures. Furthermore, Chinese does not transfer well to any other language, possibly due to differences in linguistic construction and movie distributions (See §4.5) . Lastly, we can observe that it is difficult to transfer from any language to Hindi, possibly because the pre-training data of CCLM does not contain Hindi, which hurts representation learning.
6.5 Ablation: Transfer to YMS
In this section, we extend our analysis to show that models trained on M-SyMoN can generalize to other video-text alignment benchmarks. Specifically, we evaluate baseline models trained on M-SyMoN on the out-of-domain English benchmark dataset YMS Dogan et al. (2018) and compare the results with the state-of-the-art performances.
As shown in Table 4, the large scale M-SyMoN dataset significantly improves the performance of YMS. First, we can observe that fine-tuning the base model on the human-annotated training set of YMS (i.e., CCLM-individual (YMS)) is already on par with the state-of-the-art performance, demonstrating the effectiveness of our proposed base model. Second, compared to CCLM-individual (YMS), only training on the weakly supervised portion of M-SyMoN improves performance by 13.7 and 12.7 percentage points on Clip Accuracy and Sentence IoU, respectively. This suggests the value of the large-scale weakly-supervised data in M-SyMoN. Moreover, additional fine-tuning on the human-annotated data from M-SyMoN, i.e., CCLM-two-stage-supervision (English), further boosts the performance by 2.4 and 3.0 percentage points on Clip Accuracy and Sentence IoU, respectively. These observations highlight the significance and generalizability of our human-annotated video-text correspondence dataset, indicating the potential of M-SyMoN for advancing research in video story understanding.
Clip Acc. Sent IoU
NeuMatch Dogan et al. (2018) 12.0 10.4
Wang et al. (2021a) 30.6 12.8
Sun et al. (2024) 23.2 18.4
CCLM-individual (YMS) 29.2 18.9
CCLM-multilingual 13.3 6.2
CCLM-individual (English) 35.8 25.0
CCLM-translate 25.2 16.7
CCLM-two-stage (English) 42.9 31.6
CCLM-two-stage-supervision (English) 45.3 34.6
Table 4: Out-of-domain results on the YMS dataset. CCLM-individual (YMS) is the base model fine-tuned on the human-annotated training set of YMS.
6.6 Case Study
In Figure 5, we provide a qualitative example highlighting the challenges of M-SyMoN.
First, the movie descriptions are intended to advance the story plot and do not describe the video content exactly, making alignment difficult. Second, the sentences and clips in a movie are very similar causing difficulty in alignment. For example, Figure 5 shows a war film with numerous scenes of car bombings. Without understanding the story and merely judging by the visual and textual content, the description “Vehicles set on fire during an attack” can be aligned with any of these scenes. Finally, the translation model may incur errors. For example, the last sentence in the second line of Figure 5 contains translation errors and is incomprehensible. CCLM-translate mistakenly aligns this sentence to a scene of the gunfire because “attack” appears in the sentence, while CCLM-two-stage, with access to the correct Chinese caption, makes the correct prediction.
7 Conclusion
In this paper, we introduced M-SyMoN, a large-scale multilingual video story dataset. It contains 13,166 movie summary videos from 7 languages, featuring 480 videos with human-annotated video-text correspondence. We established several multilingual multimodal baselines to benchmark M-SyMoN. Experimental results demonstrate the value of M-SyMoN for video story understanding. Its size, multilingual nature, and rich alignment annotations make M-SyMoN a valuable contribution.
Limitations
The preprocessing of M-SyMoN involves several automatic techniques that may introduce noise. First, some text descriptions are generated using Google automatic speech recognition. Automatic punctuation restoration is then applied to the text. For the translation baselines, the text is further translated into English. This pipeline may propagate errors. However, due to the size and complexity of M-SyMoN, manual processing is impractical. We acknowledge the potential for improvement in preprocessing steps as better techniques become available.
Additionally, human annotation of video-text correspondence in story videos can be ambiguous. Annotators received a list of instructions for ambiguous cases. For example, when text describes a character’s emotional state, annotators are instructed to match only when the emotion is evident from the character’s expression or action. In cases outside the provided guidelines, annotators used their best judgment. Despite the inherent ambiguity in story content annotation, manual inspection of the annotated data indicates good quality, as shown in Table 5.
Ethics Statement
M-SyMoN is constructed with YouTube videos under the Standard YouTube License. For the videos, we release a list of YouTube URL the users can use to download the video from YouTube, as is the standard practice Bain et al. (2020); Miech et al. (2019); Caba Heilbron et al. (2015). A fair compensation amount for the annotators was determined by the annotation company based on the difficulty level and time needed to annotate a minute of video in each language. On average, we compensate the annotator 2.44 USD for annotating a minute of video. The exact per-minute compensation amount for each language is shown in Table 6 in the Appendix.
In this paper, we collect user-uploaded videos from YouTube, which are summaries of mostly western movies and TV shows. We recognize that movies and TV shows are fictional in nature, and often prioritize dramatic events over faithful representation of real-life scenarios. In addition, the videos may reflect particular bias of the creators of the movie and TV shows or the creators of the summary videos, as well as bias from particular cultures or the time periods of production.
For these reasons, we urge researchers to take caution when attempting to learn social norms from such videos. For example, events of bank robberies may be over-represented in these videos, and a machine learning model may inadvertently infer that robbing a bank is part of the social norm. In addition, the model may incorrectly learn from disproportional association of certain groups of people with certain social status, occupations, and other cultural constructs. The dataset is intended for fundamental research and not real-world deployment.
Appendix A Additional Information on M-SyMoN
A.1 Additional Preprocessing
Removing Overlap Between Train and Test Sets.
In movie recap videos, videos summarizing the same movie may have similar visual and textual content. Thus, we remove all movies that appear in the test sets from the training sets. Specifically, we use ChatGPT to identify the corresponding movie for each video and translate the movie name to English. We then remove the overlapped videos based on the English movie names.
A.2 Additional Analysis
As shown in Figure 6, creators using different languages prefer similar genre. Although interestingly, “Fantasy” films are featured more in Hindi and Chinese than other language, perhaps due to the significant role mythology plays in Eastern cultures.
In Figure 7, we list the language of release for movies recapped in each language. This trend largely follows the same pattern as the country of release likely because most countries produce movies in their native language.
For each language, we list the top 5 years the movies were released (Figure 8). As our dataset collection ends around February 2022, M-SyMoN largely contains movies released from 2017 to 2021.
In Figure 9, we show the narrative styles of movies recapped in each language. The most common narrative styles are linear and episodic.
Figure 9: Narrative style of movies recapped in each language
Appendix B Annotation Details
Characteristics Of Annotators.
We employ a team of annotators with professional qualifications in translation for annotation. The team is from Korea. We discussed the compensation amount with the annotation team and set an adequate amount.
Annotation Instruction.
We gave the annotators an instruction, shown in Figure 10, and a set of annotation guidelines for ambiguous situations.
You will see a 5-15 minute video, which is a summary of a movie or a TV show. You will be given a list of sentences that describe the content of the video. Please align each sentence to the video content by writing down the start time and end time of the video segment that matches the sentence. If the sentence does not match any video content, please mark it as ‘‘Unmatched’’.
Figure 10: Human annotation instruction.
Annotation Procedure.
We divide the data randomly into 4 equal batches, each batch contains a quarter of data from each language. After the annotators annotate each batch, we randomly select 2 videos from each language for validation and employ a student familiar with the task to annotate them. Then, we calculate the IoU between the student’s annotations and the annotations provided by the annotators. Specifically, for each sentence we calculate the IoU between the duration annotated by the student and the duration annotated by the annotators. If the IoU on a particular language falls below 60%, the annotators were asked to redo the annotations for that language in that batch. The average IoU of each language is shown in Table 5
English Chinese Spanish French Portuguese Hindi Russian
IoU 83.2% 78.8% 84.4% 81.3% 88.9% 84.4% 80.8%
Table 5: IoU between annotator annotations and our annotations.
English Chinese Spanish French Portuguese Hindi Russian
Annotation Cost 1.60 1.44 2.83 3.78 2.83 1.79 2.83
Table 6: Annotator compensation amount for annotating a minute of video, in USD.
Appendix C Multilingual Punctuation Restoration
Task. We treat punctuation restoration as a token classification task, where the model predicts the appropriate punctuation, if any, to follow each token in the sequence. To simplify the task, we limit the predictions to three punctuations: period, comma, and question mark. Other punctuations are either removed (in the case of #,@," and ’) or replaced with periods and commas (in the case of !, ; and :).
Dataset.
For training, we use the Wiki-40B Guo et al. (2020) dataset containing clean Wikipedia articles in 40+ languages. For Hindi, we further supplement the training set with data from Wikipedia Hindi  [^3] https://www.tensorflow.org/datasets/catalog/wikipedia.
To evaluate our punctuation restoration model on narrative content, we build an evaluation set from Wikipedia plot summaries. Specifically, we first identify Wikipedia articles of popular movies. Then, we extract the “Plot Summary” portion of the articles. Finally, we remove any overlap between training and test data. We acquire the evaluation sets for French, Spanish, Portuguese, Chinese, and Russian following the above procedure. However, for Hindi, we can not find enough Wikipedia plot summaries. Therefore, we use the wiki-40B validation set for evaluation which may inflate performance on Hindi.
Model.
Following Frank and Böhme (2021), we fine-tune on the XLM-RoBERTa-large Conneau et al. (2019) model.
Evaluation.
We evaluate the model performance with the F1 score.
Baselines.
We use the Full-Stop Frank and Böhme (2021) model as the baseline  [^4] The original Full-Stop model is trained on English, German, French, and Italian. We follow their instruction to train on Spanish, French, and Portuguese. We use the same backbone as Full-Stop, the difference is that they train on the Europarl Koehn (2005) dataset while we train on Wiki-40B. The Europarl dataset contains transcripts of political talks in European languages and is commonly used in training multilingual punctuation restoration models.
Results.
Table 7 shows the multilingual punctuation model performance. Our model achieves good performance on all languages. On European languages, our model outperforms Full-Stop by 6-14% (see Table 8). This demonstrates that our model is suitable for punctuation restoration in narratives. The performance on Hindi is superior to other languages, this is likely because Hindi is evaluated on in-domain data from Wiki-40B.
0 . , ? Total
French 0.98 0.86 0.71 0.56 0.78
Spanish 0.97 0.77 0.67 0.53 0.74
Portuguese 0.98 0.83 0.72 0.50 0.76
Russian 0.96 0.73 0.77 0.51 0.74
Chinese 0.99 0.82 0.83 0.61 0.81
Hindi 0.99 0.94 0.88 0.77 0.90
Average 0.98 0.82 0.76 0.58 0.79
Table 7: Result of multi-lingual punctuation.
0 . , ? Total
Wiki-40B (Ours)
French 0.98 0.86 0.71 0.56 0.78
Spanish 0.97 0.77 0.67 0.53 0.74
Portuguese 0.98 0.83 0.72 0.50 0.76
Europarl
French 0.97 0.80 0.64 0.47 0.72
Spanish 0.97 0.70 0.59 0.15 0.60
Portuguese 0.97 0.75 0.63 0.37 0.68
Table 8: Comparison between training on Wiki-40B and training on Europarl. Europarl only contains European languages.
English Chinese Spanish French Portuguese Hindi Russian
Clip. Sent. Clip. Sent. Clip. Sent. Clip. Sent. Clip. Sent. Clip. Sent. Clip. Sent.
Weakly supervised
CCLM-multilingual 13.4 7.3 23.5 12.8 19.5 8.9 17.6 8.8 17.4 7.2 11.8 3.8 15.1 7.0
CCLM-individual 33.2 22.3 36.6 24.6 25.7 13.6 24.2 11.2 24.8 12.4 17.7 7.3 21.9 11.2
CCLM-translate 28.4 18.4 27.7 15.9 24.0 12.5 24.3 13.8 21.9 10.7 18.6 8.2 20.1 10.8
CCLM-two-stage 33.7 22.6 43.7 33.3 27.2 15.2 27.8 16.3 27.0 14.6 20.9 9.9 24.5 13.3
Supervised
CCLM-translate-supervised 30.9 20.0 27.5 20.1 26.5 14.0 26.9 15.6 24.4 12.3 20.9 9.3 21.9 11.8
CCLM-two-stage-supervised 34.6 23.5 45.2 34.4 27.4 15.9 28.3 16.9 27.7 15.4 20.5 9.6 25.1 13.7
Table 9: Intra-lingual results based on Clip Accuracy and Sentence IoU. Due to space limitation, we write Clip Accuracy as “Clip.” and Sentence IoU as “Sent.”
In Table 9, we show the Clip Accuracy and Sentence IoU for the intra-lingual experiments.
In Tables 10 and 11 we show the Clip Accuracy and Sentence IoU scores for cross-lingual transfer, corresponding to Figure 4
English Chinese Spanish French Portuguese Hindi Russian
CCLM-translated 28.4 27.7 24.0 24.3 21.9 18.6 20.1
  CCLM-two-stage (English) 33.7 25.4 23.3 25.9 21.1 15.1 21.4
CCLM-two-stage (Chinese) 22.4 43.7 16.7 14.8 11.8 12.0 15.3
CCLM-two-stage (Spanish) 32.5 29.9 27.2 26.6 24.2 14.8 22.0
CCLM-two-stage (French) 31.7 30.9 25.1 27.8 23.1 15.0 22.5
CCLM-two-stage (Portuguese) 32.6 30.2 25.4 27.1 27.0 16.6 22.2
CCLM-two-stage (Hindi) 31.8 29.8 25.2 26.1 21.7 20.9 22.7
CCLM-two-stage (Russian) 33.2 31.7 25.6 26.5 23.5 16.6 24.5
Table 10: Clip Accuracy scores for cross-lingual transfer.
English Chinese Spanish French Portuguese Hindi Russian
CCLM-translated 18.4 15.9 12.5 13.8 10.7 8.2 10.8
  CCLM-two-stage (English) 22.6 14.2 11.7 14.8 10.8 2.9 11.5
CCLM-two-stage (Chinese) 13.8 33.3 6.7 7.6 4.2 1.4 6.6
CCLM-two-stage (Spanish) 21.5 18.0 15.2 15.6 12.8 4.6 11.7
CCLM-two-stage (French) 20.8 17.8 13.1 16.3 11.8 3.9 15.2
CCLM-two-stage (Portuguese) 21.5 17.9 13.7 15.7 14.6 5.2 11.8
CCLM-two-stage (Hindi) 20.8 17.4 12.5 14.8 11.0 9.9 11.8
CCLM-two-stage (Russian) 22.1 19.2 13.6 15.1 12.0 5.7 13.3
Table 11: Sentence IoU scores for cross-lingual transfer.
Appendix D Video-Text Alignment
D.1 Drop-DTW
DTW uses dynamic programming to find the optimal alignment between two sequences based on distance (or similarity), the final alignment corresponds to the shortest distance or highest similarity. In the traditional DTW algorithm, each item in one sequence must match with an item in the other sequence. However, in story videos, some text are not grounded in the video and vice-versa. Therefore, we use the Drop-DTW Dvornik et al. (2021) algorithm to facilitate dropping certain clips and sentences.
In traditional DTW, to align a sequence of video clips $V=(v_{1},\ldots,v_{N})$ and a sequence of sentences $T=(t_{1},\ldots,t_{M})$, we first assume that $v_{1}$ is aligned to $t_{1}$. Thus, the cost of matching $v_{1}$ and $t_{1}$ is $c(1,1)=0$, and the cost of match $v_{1}$ with $t_{j}(j\neq 1)$ is $c(1,j)=\infty$, and vice versa. Then, the minimal cost of aligning $(v_{1},\ldots,v_{i})$ with $(t_{1},\ldots,t_{j})$, can be calculated as:
$\begin{split}c(i,j)=\min(c(i-1,j)+d(i,j),\\ c(i,j-1)+d(i,j),\\ c(i-1,j-1)+d(i,j))\end{split}$ (3)
where $d(i,j)$ denotes the distance between $v_{i}$ and $t_{j}$. Since we have the cosine similarity between each video-text pairs, the distance can be calculated as $d(i,j)=1-s(i,j)$, where $s(i,j)$ is the cosine similarity between $v_{i}$ and $t_{j}$
In Drop-DTW, we further define two hyperparameters $d_{v}$ and $d_{t}$ as the cost of dropping the video and the text respectively. At each time stamp, we calculated the minimal cost as the lowest cost between: (1) matching $v_{i}$ and $t_{j}$ and adding $d(i,j)$ to the total cost; (2) dropping $v_{i}$ and adding $d_{v}$ to the total cost; (3) dropping $t_{j}$ and adding $d_{t}$ to the cost; (4) dropping both $v_{i}$ and $t_{j}$ and adding both $d_{v}$ and $d_{t}$ to the cost.
D.2 Addition results
In Table 9 we show the Clip Accuracy and Sentence IoU scores corresponding to Table 3.
D.3 Licensing Information
The videos we acquire from YouTube are under the Standard YouTube license. Note that in M-SyMoN, the videos are released in the form of YouTube URL, and researchers can download the videos directly from YouTube.
For multilingual punctuation restoration we train with data from Wiki-40B Guo et al. (2020) licensed under the Apache 2.0 License. For scene segmentation, we use TransNet-V2 Souček and Lokoč (2020) released under the MIT license. ChatGPT, which we use for extracting movie names and metadata, is under the GNU Affero General Public License Version 3.
For video text alignment, the CCLM Zeng et al. (2023) is under the BSD-3-Clause license and the CLIP4Clip is under the MIT License. YMS Dogan et al. (2018), which we use as an out-of-domain evaluation benchmark, is from https://github.com/RubbyJ/Data-efficient-Alignment.
The usage of all model and data within the paper are in line with their intended uses.